OpenStudy (anonymous):

Am I right or wrong? If h(x)=(fog)(x) and h(x)=4 sqrt x+7, find g(x) if f(x)=4 sqrt x+1 I worked it out, but different answers a couple of times.

OpenStudy (anonymous):

i would say \(g(x)=x+6\)

OpenStudy (anonymous):

7then \[f(g(x))=f(x+6)=4\sqrt{x+6+1}=4\sqrt{x+7}\]

OpenStudy (anonymous):

@satellite73 so 7 is the answer?

OpenStudy (anonymous):

?

OpenStudy (johnweldon1993):

\[\large g(x) = x + 6\] as stated above

OpenStudy (anonymous):

maybe i am not reading the question correctly a screen shot or an equation editor would help on this one

OpenStudy (anonymous):

@satellite73 You're right, sorry. I looked at it wrong. Thank you! :)